OpenStudy (averageniick):

The sum of four consecutive odd integers is –72. Write an equation to model this situation, and find the values of the four integers.

OpenStudy (danjs):

Let say 'n' be one of the odd integers, then the next consecutive one will be 2 more, then the third 2 more again... the four consecutive odd numbers are then - n n+2 n+4 n+6

OpenStudy (danjs):

It says the sum of those 4 numbers will equal -72. That gives you.. n + n+2 + n+4 + n+6 = -72 4n + 12 = -72 4n=-84 n=-21 that is the odd integer n = -21, the others are 2,4,and6 added to that n = -21 n+2 = -19 n+4 = -17 n+6 = -15 notice those 4 consecutive odd integers add up to -72